A connical funnel of diameter 23.2cm and depth of 42cm is full of water.If the water is poured into a cylinder tin of diamter 16.2cm find the least possible height of the tin if it must contain all the liquid.
the answer at the back of the text book is 28.7
please give me a detailed answer with the steps and solution and explain how to do questions like these.

Answer by mananth(16946)

A connical funnel of diameter 23.2cm and depth of 42cm is full of water.If the water is poured into a cylinder tin of diamter 16.2cm find the least possible height of the tin if it must contain all the liquid.
A connical funnel of diameter 23.2cm , radius = 11.6 cm
and depth of 42cm = h
Volume of cone = 1/3 *pi*r^2*h
=1/3 * pi * 11.6^2*42
=1883.84pi
this is the volume of cylinder r= 8.1
1883.84pi = pi *8.1^2*h
h= 1883.84/8.1^2
h=28.71 cm

If the water is poured into a cylinder tin of diamter 16.2cm find the least possible height of the tin if it must contain all the liquid.